OpenStudy (anonymous):

Find the sum of the infinite series if it exists. If it exists, round to one decimal place. If it does not exist, write none. 4 + 3.2 + 2.56 + 2.048

OpenStudy (tkhunny):

You must find the common ratio. Let's see it.

OpenStudy (anonymous):

is it 0.8??

OpenStudy (tkhunny):

4*0.8 = 3.2 3.2*0.8 = 2.56 2.56*0.8 = 2.048 Looks like it. Why do you doubt? Now what?

OpenStudy (anonymous):

i doubt because im bad at math... but i dont know i wasnt there for this part so i didnt learn it

OpenStudy (tkhunny):

Well, stop it! You're doing fine. Whoever told you you were bad at math was just trying to discourage you. Don't listen to it. You should be learning. A little bit of a struggle is not a bad thing at all. I'm actually encouraged by it. It means you are trying and you might even care! :-) Same things as all such infinite Geometric Series: \(\dfrac{a_{0}}{1-r}\) \(a_{0}\) is hte first term \(r\) is the common ratio. If the ratio is between 1 and -1, it all converges and everything is fine!